What is the domain of (5x^2-5x+2)/(4-x^2)?
Answer by stanbon(75887)

What is the domain of (5x^2-5x+2)/(4-x^2)?
----
The denominator cannot be zero:
4-x^2 = 0
(2-x)(2+x) = 0
x = 2 or x = -2
------
The domain is All Real Numbers except x = 2 or -2.
------------
Cheers,
Stan H.
==========